# Global Compression Commander: Plug-and-Play Inference Acceleration for High-Resolution Large Vision-Language Models

### v5

@@ Abstract
Code: https://github.com/xuyang-liu16/GlobalCom2.
@@ 1 Introduction
To enhance the efficiency of LVLMs, recent efforts have increasingly adopted token compression approaches (Cha et al. 2024; Shang et al. 2025), which reduce visual tokens while preserving essential information. These architecture-agnostic methods achieve optimal efficiency-accuracy trade-offs and have proven effective for LVLM acceleration. However, these methods were primarily designed for traditional single-view architectures (the entire image in Figure 1 top). While the multi-view dynamic cropping approach enables more fine-grained visual understanding, it further increases the number of visual tokens and computational overhead.
@@
However, directly applying existing token compression methods to HR-LVLMs overlooks three critical issues: (i) Global context neglect: Current methods disregard the crucial role of global thumbnails in extracting holistic context and guiding visual understanding (Figure 2). (ii) Information richness disparity: They fail to account for varying information density across different crops (Figure 3), resulting in performance gaps exceeding 5% on high-resolution tasks when comparing the removal of most versus least informative crops. (iii) Content-agnostic positional bias: Inner-LLM question-aware compression methods (Chen et al. 2024b; Xing et al. 2025; Zhang et al. 2025) systematically allocate more tokens to later-positioned crops irrespective of their actual informative content (Figure 4), inducing severe multi-modal hallucinations under extreme compression (Table 1), with detailed analysis provided in Appendix (Figure 9). These three oversights collectively result in over-compression of information-rich regions and disruption of the visual-semantic hierarchy (Figure 7).
@@
Building on these observations, we propose “Global Compression Commander” (GlobalCom2), which follows a “global-to-local” guided token compression philosophy tailored for dynamic cropping-based HR-LVLMs. In Figure 1, GlobalCom2 leverages holistic information from thumbnails to evaluate each crop’s information richness, adaptively adjusting compression intensity. It performs token compression through comprehensive evaluation from both global and local perspectives, achieving differentiated compression across regions while preserving significant information. This training-free approach extends seamlessly to video large language models (VideoLLMs) for efficient video understanding and can be integrated with existing question-aware methods. Notably, integrating this “global-to-local” design with FastV (Chen et al. 2024b) and SparseVLM (Zhang et al. 2025) achieves significant improvements of 5.3% and 5.2% across benchmarks while alleviating multi-modal hallucination caused by positional bias.
@@
  • Superior Performance-Efficiency Trade-offs: Extensive experiments on multiple HR-LVLMs and VideoLLMs demonstrate exceptional balance of GlobalCom2, maintaining over 90% performance with 90% token reduction, while achieving substantial memory savings and throughput improvements.
@@ 2 Related Work
High-resolution LVLMs. Large vision-language models (LVLMs) integrate vision encoders and LLM decoders via projectors for feature alignment. Early LVLMs (Liu et al. 2024a; Bai et al. 2023) resize images to fixed resolutions, causing shape distortion and detail loss. To address this, high-resolution LVLMs have emerged in three categories: (i) Hybrid resolution methods using dual visual encoders (Li et al. 2024c; Luo et al. 2025). (ii) Native resolution methods with NaViT-style encoders (Wang et al. 2024; Guo et al. 2025). (iii) Dynamic cropping methods (Li et al. 2025; Zhu et al. 2025), which split images into regions for single-encoder processing. Dynamic cropping has gained widespread adoption (Lu et al. 2025; Chen et al. 2025) due to its vision encoding efficiency. However, increased visual tokens introduce computational challenges in inference speed and memory usage. Our work focuses on improving the efficiency of dynamic cropping-based HR-LVLMs.
Token Compression for LVLMs. Token compression has been widely adopted for LVLM acceleration through training-free approaches at two stages: (i) Vision Encoding stage (Shang et al. 2025; Yang et al. 2025): FasterVLM (Zhang et al. 2024) leverages [CLS] attention scores to prune visual tokens. (ii) LLM Pre-filling stage (Xing et al. 2025; Ye et al. 2025): FastV (Chen et al. 2024b) prunes tokens based on LLM self-attention, while SparseVLM (Zhang et al. 2025) uses cross-modal attention scores. Some methods combine both stages (Liu et al. 2024c). However, these methods treat all visual tokens equally in a “flat” token space, ignoring HR-LVLMs’ hierarchical structure where thumbnails and crops serve distinct roles.
@@ 3.1 Preliminary
Post-processing. LLaVA-NeXT removes padding and adds boundary tokens to preserve aspect ratios.
@@ 3.2 Stepping into Dynamic Cropping
(I) Functions of Thumbnail and Crops. We investigate how thumbnail and crops contribute to image understanding in HR-LVLMs through experiments with LLaVA-NeXT-7B.
Figure 2 shows that thumbnail alone performs better on general visual perception tasks (e.g., SQA (Lu et al. 2022) and MMBench (Liu et al. 2024d)), benefiting from holistic visual information. In contrast, local crops excel in fine-grained tasks like text understanding (TextVQA (Singh et al. 2019)) and hallucination detection (POPE (Li et al. 2023)) by providing detailed features. Therefore, we identify that:
@@
(II) Information Richness in Different Crops. We investigate visual representation variance across crops through qualitative and quantitative analysis.
Figure 5 shows upper crops contain rich content (e.g., football players) while lower crops show mainly grass. CLIP-ViT’s last-layer attention distribution confirms higher semantic information in upper regions.
Quantitative experiments measuring crop importance via [CLS] attention scores and drop-crop tests on LLaVA-NeXT-7B (Figure 3) reveal substantial performance gaps: averaging 2.4% across benchmarks and reaching 5.5% and 5.2% on high-resolution tasks MM-Vet (Yu et al. 2024) and TextVQA (Singh et al. 2019) when dropping most versus least important crops. We thus identify:
@@ 4 Global Compression Commander
HR-LVLMs with dynamic cropping increase token length, making self-attention’s quadratic complexity a bottleneck. Inspired by human vision’s process of grasping scene gist before focusing on details, we propose “Global Compression Commander” (GlobalCom2) which implements a “global-to-local” compression strategy.
In Figure 5, for thumbnail token compression, we identify tokens with essential holistic information. Given the [CLS] token’s effectiveness as global image representation (Liang et al. 2022), GlobalCom2 uses the last ViT layer’s attention map to compute attention between each token and [CLS] [^1] For HR-LVLMs without [CLS] token, we propose an alternative token informativeness evaluation metric. See Appendix. (blue path in Figure 5). For the 1D token sequence $\mathbf{X}^{G}$ of length $N$, the importance score ${s}_{i}^{G}$ for the $i$-th token is:
@@ 4.2 Holistic Token Evaluation
where we empirically set $\alpha=0.5$ to give equal consideration to both information sources (ablation study of $\alpha$ is in Appendix Figure 10). The final compression:
@@
4.3 Extension to Video Understanding
5 Experiments
Method MVBench LongVideoBench MLVU VideoMME Average
Overall Short Medium Long
Upper Bound, 6272 Tokens
LLaVA-OneVision-7B 56.9 56.4 63.0 58.6 70.3 56.6 48.8 100.0%
Ratio=30%, Retain 1882 Tokens
DyCoke(CVPR25) 56.6 54.7 60.3 56.1 67.1 54.6 46.6 96.5%
Ratio=25%, Retain 1568 Tokens
FastV(ECCV24) 55.5 53.3 59.6 55.3 65.0 53.8 47.0 95.0%
PDrop(CVPR25) 55.3 51.3 57.1 55.5 64.7 53.1 48.7 94.2%
SparseVLM(ICML25) 56.4 53.9 60.7 57.3 68.4 55.2 48.1 97.5%
FrameFusion(ICCV25) 56.0 54.8 61.7 57.5 68.2 55.7 48.6 98.1%
GlobalCom2 57.0 55.4 62.6 58.1 69.2 55.8 49.4 99.3%
Ratio=15%, Retain 941 Tokens
FastV(ECCV24) 51.6 48.3 55.0 48.1 51.4 49.4 43.3 85.0%
PDrop(CVPR25) 53.2 47.6 54.7 50.1 58.7 48.7 45.0 87.4%
SparseVLM(ICML25) 52.9 49.7 57.4 53.4 61.0 52.1 47.0 91.2%
GlobalCom2 54.2 52.9 60.7 55.8 66.0 53.3 48.1 95.3%
5.1 Experimental Setting
We evaluate on LLaVA-NeXT (Liu et al. 2024b) and LLaVA-OneVision (Li et al. 2025) for image understanding, LLaVA-OneVision and Qwen2-VL (Wang et al. 2024) for video understanding. We compare with FastV (Chen et al. 2024b), SparseVLM (Zhang et al. 2025), PDrop (Xing et al. 2025), PruMerge (Shang et al. 2025), FasterVLM (Zhang et al. 2024), and VideoLLM methods DyCoke (Tao et al. 2025) and FrameFusion (Fu et al. 2025) at different retention ratios $R$. For fair comparison with multi-stage methods, we use “equivalent token count” reflecting the average percentage of vision tokens retained across all LLM layers.
5.2 Main Results
Method LongVideo. MLVU VideoMME
Overall Short Medium Long
Upper Bound
Qwen2-VL-7B 56.0 60.4 57.6 70.0 54.6 48.2
Ratio=25%
FastV(ECCV24) Out of Memory (OOM)
DyCoke(CVPR25) 50.5 55.2 51.5 62.1 47.1 45.1
GlobalCom2 51.4 55.9 54.3 64.7 50.1 48.2
@@ 5.2 Main Results
Image Understanding. Table 1 compares GlobalCom2 with existing methods with LLaVA-NeXT, which demonstrates its three key advantages: (i) Superior performance: GlobalCom2 consistently outperforms all baselines, being the only method that maintains >90% of the original performance across all retention ratios. (ii) Extreme compression robustness: While baseline methods suffer significant degradation at $R=10\%$, GlobalCom2 maintains robust performance and achieves the best results across all benchmarks, outperforming the second-best method by 1.7% on average. Notably, FastV and PDrop demonstrate severe deterioration on POPE, exhibiting clear multi-modal hallucination due to positional bias from attention-guided token selection (Figure 7). (iii) High-resolution excellence: GlobalCom2 demonstrates exceptional performance on high-resolution benchmarks (e.g., VQAT, MME, MM-Vet). Our “global-to-local” guided compression preserves both global semantics and local details, outperforming baselines that suffer from over-compression. Figure 6 further validates GlobalCom2’s effectiveness on LLaVA-OneVision, maintaining 90.5% of the original performance at $R=10\%$ while consuming only 35.4% of the original GPU memory. See Appendix for results on the larger HR-LVLM.
Video Understanding. Table 2 presents GlobalCom2’s extension to video understanding with LLaVA-OneVision, demonstrating two key advantages: (i) Superior VideoLLM performance: GlobalCom2 achieves 2.7% average improvement over VideoLLM-specific method DyCoke (Tao et al. 2025) while using fewer vision tokens. (ii) Advantages on long video understanding: At $R=15\%$, GlobalCom2 achieves 3.2 and 3.3 points higher than the second-best method on LongVideoBench and MLVU, demonstrating the effectiveness of our “global-to-local” guided compression design. Table 3 shows GlobalCom2 significantly outperforms DyCoke on Qwen2-VL, confirming its scalability. Additionally, FastV explicitly computes the full attention matrix in the certain layer, causing OOM issues with high token counts and limiting its practical applications.
@@ 5.4 Efficiency Analysis
Table 6 compares inference efficiency among SparseVLM, FasterVLM and GlobalCom2. Unlike SparseVLM which requires explicit attention scores in LLM and is naturally incompatible with FlashAttention (Dao et al. 2022), FasterVLM and our GlobalCom2 enable efficient computation before LLM decoding for efficient computation. As a plug-and-play solution, GlobalCom2 achieves superior performance-efficiency trade-offs, maintaining 90% of the original performance while dramatically reducing peak memory usage by 40% and boosting inference throughput by $1.8\times$. Detailed analysis is in Appendix Table 10.
@@ 5.5 Compression Visualizations
Figure 7 compares compression processes of different methods at $R=10\%$ on VQAT, revealing over-compression issues in baselines: (i) Positional Bias: FastV exhibits clear positional bias, allocating more tokens ($3\times$) to later-positioned crops regardless of visual content. (ii) Uniform Compression: FasterVLM applies uniform compression across all crops, treating them as equally important. It fails to preserve critical information in some regions while retaining redundancy in others. In contrast, GlobalCom2 globally assesses crop informativeness and adaptively adjusts compression ratios, preserving crucial details while eliminating redundancy. See Appendix Figure 11 for more examples.
Method FLOPs↓ Memory↓ Throughput↑ Performance↑
(T) (GB) (samples/s)

### v6

@@ Abstract
Codehttps://github.com/xuyang-liu16/GlobalCom2
Extended versionhttps://arxiv.org/abs/2501.05179
@@ 1 Introduction
To enhance the efficiency of LVLMs, recent efforts have increasingly adopted token compression approaches (Cha et al. 2024; Shang et al. 2025; Liu et al. 2025b, c, a; Wen et al. 2025c), which reduce visual tokens while preserving essential information. These architecture-agnostic methods achieve optimal efficiency-accuracy trade-offs and have proven effective for LVLM acceleration. However, these methods were primarily designed for traditional single-view architectures (the entire image in Figure 1 top). While the multi-view dynamic cropping approach enables more fine-grained visual understanding, it further increases the number of visual tokens and computational overhead.
@@
However, directly applying existing token compression methods to HR-LVLMs overlooks three critical issues: (i) Global context neglect: Current methods disregard the crucial role of global thumbnails in extracting holistic context and guiding visual understanding (Figure 2). (ii) Information richness disparity: They fail to account for varying information density across different crops (Figure 3), resulting in performance gaps exceeding 5% on high-resolution tasks when comparing the removal of most versus least informative crops. (iii) Content-agnostic positional bias: Inner-LLM question-aware compression methods (Chen et al. 2024a; Xing et al. 2025; Zhang et al. 2025) systematically allocate more tokens to later-positioned crops irrespective of their actual informative content (Figure 4), inducing severe multi-modal hallucinations under extreme compression (Table 1). These three oversights collectively result in over-compression of information-rich regions and disruption of the visual-semantic hierarchy (Figure 7).
@@
Building on these observations, we propose “Global Compression Commander” (GlobalCom2), which follows a “global-to-local” guided token compression philosophy tailored for dynamic cropping-based HR-LVLMs. In Figure 1, GlobalCom2 leverages holistic information from thumbnails to evaluate each crop’s information richness, adaptively adjusting compression intensity. It performs token compression through comprehensive evaluation from both global and local perspectives, achieving differentiated compression across regions while preserving significant information. This approach can be integrated with existing question-aware methods. Notably, integrating this “global-to-local” design with FastV (Chen et al. 2024a) and SparseVLM (Zhang et al. 2025) achieves significant improvements of 5.3% and 5.2% across benchmarks while alleviating multi-modal hallucination caused by positional bias.
@@
  • Superior Performance-Efficiency Trade-offs: Extensive experiments on multiple HR-LVLMs with dynamic cropping demonstrate exceptional balance of GlobalCom2, maintaining over 90% performance with 90% token reduction, while achieving substantial memory savings and throughput improvements.
@@ 2 Related Work
High-resolution LVLMs. LVLMs integrate vision encoders and LLM decoders via projectors for feature alignment (Liu et al. 2023; Dai et al. 2023). Early LVLMs (Liu et al. 2024a; Bai et al. 2023) resize images to fixed resolutions, causing shape distortion and detail loss. To address this, high-resolution LVLMs have emerged in three categories: (i) Hybrid resolution methods using dual visual encoders (Li et al. 2024c; Luo et al. 2025). (ii) Native resolution methods with NaViT-style encoders (Wang et al. 2024; Guo et al. 2025). (iii) Dynamic cropping methods (Li et al. 2025; Zhu et al. 2025), which split images into regions for single-encoder processing. Dynamic cropping has gained widespread adoption (Lu et al. 2025; Chen et al. 2025) due to its vision encoding efficiency. However, increased visual tokens introduce computational challenges in inference speed and memory usage. Our work focuses on improving the efficiency of dynamic cropping-based HR-LVLMs.
Token Compression for LVLMs. Token compression, which aims to reduce the sequence length of tokens for computation efficiency, has been widely adopted for model acceleration (Rao et al. 2021; Liang et al. 2022; Bolya et al. 2023). For LVLMs, recent works have focused on training-free token compression for LVLM acceleration at two stages: (i) Vision Encoding stage (Shang et al. 2025; Yang et al. 2025): FasterVLM (Zhang et al. 2024) leverages [CLS] attention scores to prune visual tokens. (ii) LLM Pre-filling stage (Xing et al. 2025; Ye et al. 2025; Wen et al. 2025b): FastV (Chen et al. 2024a) prunes tokens based on LLM self-attention, while SparseVLM (Zhang et al. 2025) uses cross-modal attention scores. Some methods combine both stages (Liu et al. 2024c; Han et al. 2024). However, these methods treat all visual tokens equally in a “flat” token space, ignoring HR-LVLMs’ hierarchical structure where thumbnails and crops serve distinct roles.
@@ 3.1 Preliminary
Post-processing. LLaVA-NeXT removes padding tokens and adds boundary tokens to mark image regions, preserving aspect ratios while enhancing efficiency.
@@ 3.2 Stepping into Dynamic Cropping
(I) Functions of Thumbnail and Crops: We begin with an exploratory experiment to investigate how thumbnail and crops contribute to image understanding in HR-LVLMs by using them separately as input to LLaVA-NeXT-7B.
Figure 2 demonstrates that using global thumbnail alone yields superior performance on general visual perception benchmarks (e.g., SQA (Lu et al. 2022) and MMBench (Liu et al. 2024d)). This advantage stems from global thumbnails providing holistic visual information through complete ViT encoding. In contrast, using only local crops, where ViT independently encodes each crop, shows inferior performance on these general visual perception tasks. However, local crops excel in fine-grained perception tasks like TextVQA (Singh et al. 2019) (VQAT, testing text-centric visual understanding) and POPE (Li et al. 2023) (testing model hallucination) by providing detailed visual features. Therefore, we identify that:
@@
(II) Information Richness in Different Crops: Given that dynamic cropping leads to distinct visual representations across crops, we investigate the variance through both qualitative and quantitative perspectives.
Through visual inspection of the input image in Figure 5, we observe that the upper crops contain rich visual content (e.g., football players), while the lower crops mainly show redundant grass areas. To quantitatively validate this observation, we analyze the attention scores between [CLS] and patch tokens, which have been shown to effectively indicate token importance in prior works (Liang et al. 2022; Han et al. 2024). As visualized in Figure 5, the attention distribution from CLIP-ViT’s last layer shows significantly higher values in the upper regions, confirming that upper crops contain more semantic information from a global perspective.
Building upon this qualitative analysis, we conduct quantitative studies to validate our observations. Since attention scores with the [CLS] token effectively measure token-level semantic richness, we leverage these scores to assess crop importance by by computing the sum of partitioned attention scores within each crop region. To examine crop contributions in HR-LVLMs, we conduct experiments on LLaVA-NeXT-7B by selectively dropping local crops (i.e., dropping the most/least important one). As shown in Figure 3, dropping the most versus least important crops leads to significant performance gaps across visual understanding tasks, with an average drop of 2.4% across six benchmarks and particularly notable in VQAT (5.2% gap). Based on the above analysis, we identify that:
@@ 4 Global Compression Commander
HR-LVLMs with dynamic cropping increase token length ( in LLaVA-NeXT (Liu et al. 2024b), 10× in LLaVA-OV (Li et al. 2025)), making self-attention’s quadratic complexity a bottleneck. Inspired by human vision’s process of grasping scene gist before focusing on details, we propose “Global Compression Commander” (GlobalCom2) which implements a “global-to-local” compression strategy.
In Figure 5, for thumbnail token compression, we identify tokens with essential holistic information. Given the [CLS] token’s effectiveness as global image representation (Liang et al. 2022), GlobalCom2 uses the last ViT layer’s attention map to compute attention between each token and [CLS] (blue path in Figure 5). For the 1D token sequence $\mathbf{X}^{G}$ of length $N$, the importance score ${s}_{i}^{G}$ for the $i$-th token is:
@@ 4.2 Holistic Token Evaluation
where we empirically set $\alpha=0.5$ to give equal consideration to both information sources. The final compression:
@@
4.3 GlobalCom2 without [CLS] Token
For HR-LVLMs without [CLS] token (e.g., LLaVA-OneVision (Li et al. 2025) with SigLIP (Zhai et al. 2023)), we propose an alternative token informativeness measure for GlobalCom2. Specifically, given a sequence of vision tokens $\mathbf{X}\in\mathbb{R}^{N\times d}$ after vision encoding, we first compute a global mean vector $\mathbf{g}\in\mathbb{R}^{d}$ through global average pooling over all tokens, and then calculate the cosine similarity between each patch token $\mathbf{x}_{i}$ and $\mathbf{g}$:
$c_{i}=\cos(\mathbf{x}_{i},\mathbf{g})=\frac{\mathbf{x}_{i}\cdot\mathbf{g}}{\|\mathbf{x}_{i}\|\|\mathbf{g}\|},$ (7)
The informativeness score $s_{i}=-c_{i}$ is negatively correlated with the similarity, reflecting that tokens with greater difference from $\mathbf{g}$ carry more unique information. Specifically, tokens exhibiting low similarity to $\mathbf{g}$ represent distinctive and irreplaceable visual elements, while highly similar tokens typically correspond to redundant or common patterns. This scoring mechanism serves as an effective alternative to [CLS]-based scoring in Equation 2-5 for models without [CLS] token. It is adopted to evaluate both crop-level information richness and token-level importance in LLaVA-OneVision. We conduct comprehensive quantitative and qualitative analyses in Appendix to explore how to measure token informativeness without [CLS] token.
5 Experiments
5.1 Experimental Setting
We evaluate on LLaVA-NeXT (Liu et al. 2024b) and LLaVA-OneVision (Li et al. 2025) for evaluation. We compare with FastV (Chen et al. 2024a), SparseVLM (Zhang et al. 2025), PDrop (Xing et al. 2025), PruMerge (Shang et al. 2025), FasterVLM (Zhang et al. 2024) at different retention ratios $R$. For fair comparison with multi-stage methods, we use “equivalent token count” reflecting the average percentage of vision tokens retained across all LLM layers.
5.2 Main Results
@@
Results on LLaVA-NeXT. Table 1 compares GlobalCom2 with existing methods with LLaVA-NeXT, which demonstrates its three key advantages: (i) Superior performance: GlobalCom2 consistently outperforms all baselines, being the only method that maintains >90% of the original performance across all retention ratios. (ii) Extreme compression robustness: While baseline methods suffer significant degradation at $R=10\%$, GlobalCom2 maintains robust performance and achieves the best results across all benchmarks, outperforming the second-best method by 1.7% on average. Notably, FastV and PDrop demonstrate severe deterioration on POPE, exhibiting clear multi-modal hallucination due to positional bias from attention-guided token selection (Figure 7). (iii) High-resolution excellence: GlobalCom2 demonstrates exceptional performance on high-resolution benchmarks (e.g., VQAT, MME, MM-Vet). Our “global-to-local” guided compression preserves both global semantics and local details, outperforming baselines that suffer from over-compression.
Results on LLaVA-OneVision. Figure 6 further presents GlobalCom2’s performance across benchmarks under varying retention ratios $R$ on LLaVA-OneVision. Generally, model performance correlates with $R$, with more aggressive compression leading to degradation. Vision-centric tasks (e.g., GQA, VizWiz, MME, SEED) show substantial drops with reduced visual tokens, while SQA maintains robust performance even with minimal tokens, suggesting language understanding dominates in scientific reasoning. Notably, GlobalCom2 preserves 90.5% performance at $R=10\%$ while consuming only 35.4% of the original GPU memory, all without any training overhead, demonstrating its effective and efficient token compression.
5.3 Ablation Study and Analysis
@@ 5.4 Efficiency Analysis
Table 4 compares inference efficiency among SparseVLM, FasterVLM and GlobalCom2. SparseVLM requires explicit attention scores in LLM and is naturally incompatible with FlashAttention (Dao et al. 2022), leading to higher memory costs (Wen et al. 2025a). Instead, FasterVLM and our GlobalCom2 enable efficient computation before LLM decoding for efficient computation. As a plug-and-play solution, GlobalCom2 achieves superior performance-efficiency trade-offs, maintaining 90% of the original performance while dramatically reducing peak memory usage by 40% and boosting inference throughput by $1.8\times$.
@@ 5.5 Compression Visualizations
Figure 7 compares compression processes of different methods at extreme compression setting of $R=10\%$ on VQAT, revealing over-compression issues in baselines: (i) Positional Bias: FastV exhibits clear positional bias, allocating more tokens ($3\times$) to later-positioned crops regardless of visual content. (ii) Uniform Compression: FasterVLM applies uniform compression across all crops, treating them as equally important. It fails to preserve critical information in some regions while retaining redundancy in others. In contrast, GlobalCom2 globally assesses crop informativeness and adaptively adjusts compression ratios, preserving crucial details while eliminating redundancy.
Method TFLOPs↓ Memory↓ Throughput↑ Performance↑